OpenStudy (anonymous):

–2(6 – 20n) = 20(2n – 3) A. 40 B. –36 C. no solution D. all real numbers so i got –2(6 – 20n) = 20(2n – 3) 12-40n=40n-60 n=72

OpenStudy (igreen):

First step, distribute -2 and 20 into the parentheses.

OpenStudy (igreen):

\(-2(6 - 20n)\) Distribute -2. Do you know how to do that?

OpenStudy (anonymous):

2*6 =12

OpenStudy (igreen):

That's -2, not 2.

OpenStudy (igreen):

-2 * 6 = ?

OpenStudy (igreen):

-2 * -20n = ?

OpenStudy (lunymoony):

–2(6 – 20n) = 20(2n – 3) -12 + 40n = 40n -60 ........when two minus signs meet they turn positive

OpenStudy (anonymous):

2*6 is 12 cuz ur multipling

OpenStudy (lunymoony):

yep!

OpenStudy (anonymous):

ok

OpenStudy (lunymoony):

so "-12+40n = 40n -20

OpenStudy (lunymoony):

Have any questions just ask

OpenStudy (anonymous):

so it will be 48

OpenStudy (lunymoony):

-12+40n = 40n -20 take 40n from both sides, remember each are positive -12 = -20

OpenStudy (lunymoony):

-12+40n = 40n -60 ....sry, meant it was a 3, so it is 60

OpenStudy (lunymoony):

we now have -12 = -60

OpenStudy (anonymous):

my quieston is what do i do with the 40

OpenStudy (lunymoony):

Add 12 to each side to get a one side solution -60 (+12) 48

OpenStudy (lunymoony):

the 40n?

OpenStudy (lunymoony):

We took 40n from both sides

OpenStudy (anonymous):

ok